OpenStudy (anonymous):

Write an equation in point-slope form for the line through the given point with the given slope. (8, 3); m = 6

OpenStudy (anonymous):

Whats The Answer Choices

OpenStudy (anonymous):

y – 3 = 6(x – 8)

OpenStudy (anonymous):

A. y + 3 = 6(x - 8) B. y - 3 = 6(x - 8) C. y - 3 = 6(x + 8) D. y + 3 = 6x + 8

OpenStudy (anonymous):

so c ??

OpenStudy (anonymous):

No B

OpenStudy (anonymous):

oo yeah Word thanks :P

OpenStudy (anonymous):

Youu Welcomme ,